OpenStudy (anonymous):

Can someone please help me with this question? Medal and Fan. Find the sum of a finite geometric sequence from n = 1 to n = 5, using the expression -3(4)n - 1

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello!

OpenStudy (misty1212):

that is not such a large number of terms, you can forget the formula and add them up

OpenStudy (misty1212):

but it is not so clear what you meant it is \[-3\times 4^{n-1}\]?

OpenStudy (anonymous):

Yes, That is what I meant, I am sorry!

OpenStudy (anonymous):

-1,023 1,223 1,023 -4,374 These are my answer choices as well.

OpenStudy (misty1212):

lol use the equation tool, saves on confusion q

OpenStudy (misty1212):

just add

OpenStudy (misty1212):

\[-3\times 4^0-3\times 4^2-3\times 4^2-3\times 4^3-3\times 4^4\]

OpenStudy (misty1212):

or if you prefer \[-3(1+4+4^2+4^3+4^4)\]

OpenStudy (anonymous):

Okay, let me try and figure this out really quick.

OpenStudy (anonymous):

Do I distribute the -3 into each?

OpenStudy (anonymous):

-1023? @misty1212 :)

OpenStudy (misty1212):

idk i didn't do it hold on

OpenStudy (misty1212):

you are right \[\color\magenta\heartsuit\]

OpenStudy (anonymous):

Yay!!! Thank you so much for the help. Your way made it a lot easier! Thank you so much for the help!